In the sequence 4, 8, 64,..., each term after the first is the product of all the positive integer factors of the previous term. What is the fourth term of this sequence?
a) 2^36
b) 2^21
c) 2^12
d) 2^11

Answer by stanbon(75887)

In the sequence 4, 8, 64,..., each term after the first is the product of all the positive integer factors of the previous term. What is the fourth term of this sequence?
----
Sequence:
1st term: 2^2 = 4
2nd term: 2^1*2^2 = 2^3 = 8
3rd term: (2^1*2^2)*2^3 = 2^6 = 64
4th term: (2^1*2^2*2^3)*2^4*2^5*2^6 = 2^21
=============
Cheers,
Stan H.
==============
a) 2^36
b) 2^21
c) 2^12
d) 2^11